Company ABg offers hurricane insurance to homeowners in Texas. Assume that the premium collected for hurricane insurance can be modeled by an Ex- ponential distribution with a mean of 3. Suppose that the hurricane losses can be modeled by an Exponential distribution with a mean of 1. If the premium collected and hurricane losses are independent, what is the probability that the ratio of losses to premium collectegh will be greater than 1? A) 0.25 B) 0.28 C) 0.63 (D)p.72 E) 0.75 The circled answer is wrong please show steps to arrive at a correct answer.
Answer #1

25)

X = loss

Y = premium

X follow exponential with mean 1

Y follow exponential with mean 3

f(x,y) = f(z) * f(y) = e-r-e-3

P(X/Y > 1)

= P(X >Y)

= -r-3drdy (Decimal : 0.25) = -

option A) is correct

Please rate